# Light-Weight Diffusion Multiplier and Uncertainty Quantification for Fourier Neural Operators

### v1

@@ 4.1 Results: deterministic performance
Table 2: Evaluation of the uncertainty output quality of DINOZAUR and baseline models. The best result is bold, second best is $\underline{\text{underlined}}$
@@
GINOD $\underline{7.38}$ 18.58 0.36 16.64
GINOL 10.38 $\underline{3.15}$ $\underline{0.27}$ 15.30
TGINOD 9.18 54.13 0.41 23.17
TGINOL 9.28 3.07 $\underline{0.27}$ $\underline{14.17}$
DINOZAURD (Ours) 6.92 16.00 0.36 16.05
DINOZAURL (Ours) 8.82 $\underline{3.15}$ $\underline{0.27}$ 15.30
DINOZAURB (Ours) 7.46 4.35 0.20 12.73
@@ 4.2 Results: uncertainty quantification
In this section, we assess the UQ performance of the Bayesian version of our model, DINOZAURB, introduced in Section 3.3, on the more challenging non-uniform data. As baselines, we consider MC Dropout Gal and Ghahramani [2016] (NOD) and Laplace approximation Magnani et al. [2022] (NOL). MC Dropout is a widely used, model-agnostic UQ method, which we apply by enabling 10% dropout in the Fourier blocks’ MLPs to emulate stochasticity in the message passing. The Laplace approximation follows Magnani et al. [2022], modeling uncertainty via a fitted distribution over the parameters of the final MLP layer $\mathcal{Q}$. The discussion of hyperparameter choice, including $\mu_{\text{prior}},\sigma_{\text{prior}}$ required for setting $\ln\tau$ distributions is in the Appendix B.
UQ results in Table 2 are obtained by estimating the mean and standard deviation from 100 posterior predictive distribution samples per test data instance. Metrics are computed using the Uncertainty Toolbox Chung et al. [2021]; full metric definitions are provided in the Appendix D. DINOZAURB matches or surpasses all baselines, achieving substantially lower negative log-likelihood (NLL) than dropout models and comparable results to Laplace approximation. It outperforms all baselines in terms of miscalibration area (MA) and interval score (IS). An important consideration is that the Laplace approximation requires a secondary computationally expensive step to obtain the Hessian of the loss function with respect to the network parameters after training. In contrast, DINOZAURB natively supports UQ, thereby providing accurate uncertainty estimates out of the box. See results in Figure 3.
@@ 4.3 Ablation studies
To evaluate the scalability of DINOZAUR, we conduct experiments varying model width, depth, and the number of truncation modes using the ShapeNet car dataset. As shown in Figure 4(a), test error (measured by RL2 loss) decreases with increasing hyperparameters, saturating near the values used in our main experiments. Further improvements may be limited by the input data resolution of $64^{3}$. Figure 4(b) confirms our scaling claims: depth increases parameter count almost linearly, width affects it quadratically, while modes do not affect the model size.
@@ 5 Discussion
In our investigations, we found that the primary limitation of the diffusion mechanism is its radial symmetry. We observed that diffusion alone struggles to train well. Our solution is to include gradient features, though at the cost of increased computational time. Investigating other strategies, like introducing anisotropy intrinsically, using the modifications proposed in Liu-Schiaffini et al. [2024], or mixing gradients with other features before applying inverse FFT overhead, are promising directions for future work. We found that our Bayesian formulation is sensitive to hyperparameter settings of the prior distributions. Customizing priors across network layers to capture different physical properties could improve the robustness and interpretability of the model. More broadly, exploring other parsimonious kernel parameterizations may offer both interpretability and improved generalization, especially in settings with limited data or strong physical priors. Finally, since our parameterization is independent of spatial dimensionality, it opens up opportunities for pretraining on lower-dimensional or simpler physical systems, with the potential to accelerate learning in more complex, higher-dimensional domains through transfer learning Hao et al. [2024].
@@ Appendix A Benchmarks
Darcy flow 2 Regular [241;241][241;241][ 241 ; 241 ] Regular [241;241][241;241][ 241 ; 241 ] 1024, 1024
NS 2D+T 2+1 Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] Regular [64;64;10][64;64;10][ 64 ; 64 ; 10 ] 1000, 200
ShapeNet car 3 Regular [64;64;64][64;64;64][ 64 ; 64 ; 64 ] Irregular [3,586][3{,}586][ 3 , 586 ] 561, 100
[…]
@@
Darcy flow Permeability Standard [271;271][271;271][ 271 ; 271 ] Pressure Standard
@@
ShapeNet car SDF None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
Ahmed bodies SDF, Inlet velocity None [72;72;72][72;72;72][ 72 ; 72 ; 72 ] Pressure Standard
@@
This dataset contains steady-state 2D solutions of the Darcy flow equation on the unit box, modeling the flow of fluid through a porous medium Li et al. [2021]. It includes samples with spatial resolution of $\times 241$ 241 points. The network received a scalar permeability field as input and was tasked to predict scalar pressure values. To ensure non-periodicity, we applied padding during training, increasing the resolution to $\times 271$ 271. A standard scaler was applied to both input and output fields. Train and test splits had 1024 samples each.
@@
The Navier–Stokes dataset includes simulations of incompressible, viscous fluid flow with constant density and viscosity $\nu=10^{-5}$ Li et al. [2021]. It features a periodic 2D domain of resolution $\times 64$ 64, with 20 time steps of the vorticity field evolution. We trained models to predict the final 10 time steps from the initial 10, using 1000 training and 200 test samples. Inputs were composed as 3D tensors containing 10 time steps for every spatial point. Padding was introduced only to the time dimension, making the input to be of 64×64×1564$\times$ 64$\times$ 1564 × 64 × 15. We appended spatial coordinates uniformly sampled in $[0,2\pi)$ and a time coordinate with an integer index $t=[0,1,\dots 9]$, which resulted in 4 input channels that were mapped to one channel.
@@
This benchmark contains 661 car meshes paired with Reynolds-Averaged Navier–Stokes (RANS) simulations Umetani and Bickel [2018]. We focus on predicting the scalar pressure field at the mesh vertices, which are fixed in size at 3,586 points per mesh. The dataset is split into 561 training and 100 test samples. Target pressures were standardized during training. Following Li et al. [2024], inputs are constructed by sampling a fixed-resolution $64^{3}$ grid over the global bounding box encompassing all meshes. A signed distance function (SDF) is computed per mesh. Grids are padded to $72^{3}$ before being passed to the network.
@@
The Ahmed bodies dataset Ahmed et al. [1984] follows a similar setup to the ShapeNet car but introduces varying inlet velocities as an additional signal. This constant is appended to the SDF input, resulting in two input channels. Meshes vary in size, containing between 90,000 and 200,000 vertices. The grid resolution, padding, and target scaling match those used in the ShapeNet car.
The GINO-decoder architecture Li et al. [2024] used for both ShapeNet and Ahmed datasets requires precomputed fixed-radius neighborhoods to map grid samples to mesh vertices. We use the radii from the original paper: 0.050.050.05 for the ShapeNet car and 0.0350.0350.035 for Ahmed bodies.
For models trained on non-uniform data, we additionally apply Transformer-style positional encoding Vaswani et al. [2017] to the input features, grid locations, and mesh vertices.
@@ Appendix B Configurations of trained architectures
TGINOD 64 [24, 24, 24] $10^{-4}$ False $\mathrm{D}=0.1$
@@
TGINOL 64 [24, 24, 24] $10^{-4}$ False $C=10^{6},$
DINOZAURL 64 [24, 24, 24] No True $\alpha=500$
@@
All results reported in our paper are from experiments conducted specifically for this study to ensure a fair and consistent comparison across models. Architectural settings for the baselines are summarized in Table 5. All models featured 4 FNO blocks with GELU activations. The channel width $d_{c}$ and the number of modes were selected to align with the original configurations of FNO Li et al. [2021] and GINO Li et al. [2024], with the exception of Navier-Stokes data. We increased modes and channels to reflect the complexity of this benchmark. For models employing Tucker factorization, the rank $r$ was either adopted directly from the original implementation (GINO) or chosen to match the total number of parameters in DINOZAUR.
@@
where the expectation is approximated via Monte Carlo samples, and $C$ and $\alpha$ are hyperparameters selected on a validation set. In our experiments, we set $C=10^{6},\,\alpha=500$
@@
Optimization hyperparameters are detailed in Table 6. All models were trained using either AdamW Schedule-Free Defazio et al. [2024] or AdamW combined with a One-Cycle scheduler Smith and Topin [2019], weight decay set to $10^{-4}$. Both options allow for faster convergence with fewer epochs and reduce sensitivity to hyperparameter choices. The final setup was selected by comparing both options and reporting the results for the configuration that achieved the best performance. A limited learning rate sweep was conducted in the range $[10^{-4},10^{-2}]$, with early termination applied to underperforming runs. For the One-Cycle scheduler, the learning rate in the table refers to the max LR parameter.
@@ Appendix C Distributions of diffusion times
Figure 7: Lower diagonal full-rank learned covariance matrices of the posterior distributions. Each matrix represents a $d_{c}\times d_{c}$ ($\times 64$ 64) covariance matrix of the multi-dimensional distribution of diffusion times within each block.
@@ Appendix E Elaboration on the universal approximation
$\mathcal{N}_{\theta_{i},\tau_{i}}[v](x):=\sigma\left(W_{i}^{\mathrm{skip}}v(x)+b_{i}+W_{i}^{\mathrm{mix}}\begin{bmatrix}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}(x)\\ \mathcal{G}_{W_{i}^{\mathrm{grad}}}\Bigl{[}\mathcal{F}^{-1}\bigl{[}\exp(-4\pi^{2}\|k\|^{2}\tau_{i})\odot\mathcal{F}[v](k)\bigr{]}\Bigr{]}(x)\end{bmatrix}\right).$
@@
making a full network $\mathcal{N}_{\theta,\infty}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\infty}\circ\mathcal{P}$ an averaging neural operator in the sense of Equation 2.3 in Lanthaler et al. [2023] on the torus $\mathbb{T}^{d}$. By Corollary 3.3 Lanthaler et al. [2023], $\mathcal{N}_{\theta,\infty}$ is a universal approximator on the torus. Here, we note that the lifting and projection operators defined in Lanthaler et al. [2023] have an explicit positional dependency, and thus, one should understand them as periodized MLPs. 111 Note that Theorem 2.1 in Lanthaler et al. [2023] as it is formulated, applies to bounded Lipschitz domains. Thus, the proof of Corollary 3.3 Lanthaler et al. [2023] is not immediate. One must either follow the proof of Theorem 2.1 to see that it applies to the torus, or first formulate it on the domain $(0,2\pi)^{d}$, invoke the Theorem 2.1, and then argue that it holds when one restricts to periodic functions on $(0,2\pi)^{d}$.
Now consider a general domain $\Omega\subset(0,2\pi)^{d}$. By Lemma 41 by Kovachki et al. [2021] and Lemma A1 in Lanthaler et al. [2023], there exists a continuous, linear extension operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ and $\mathcal{E}[a]$ is periodic on $[0,2\pi]^{d}$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$. We note that the conditions of the Theorem 9 in Kovachki et al. [2021] are satisfied. By following the proof of this theorem, we apply the extension operator $\mathcal{E}$ and use the universal approximation on the torus, establishing universal approximation on $\Omega$. ∎

### v2

@@ 4.1 Results: deterministic performance
Table 2: Evaluation of the uncertainty output quality of DINOZAUR and baseline models. Mean and standard deviation values are computed from three trained models. The best result is bold, second best is $\underline{\text{underlined}}$
@@
GINOD $7.94\pm 0.54$ $3.090\pm 0.044$ $0.288\pm 0.010$ $14.731\pm 0.625$
GINOL $7.90\pm 0.29$ $3.106\pm 0.023$ $0.283\pm 0.007$ $14.709\pm 0.343$
TGINOD $7.18\pm 0.17$ $2.996\pm 0.010$ $0.278\pm 0.005$ $13.704\pm 0.189$
TGINOL $7.51\pm 0.73$ $3.071\pm 0.059$ $\underline{0.266\pm 0.018}$ $14.240\pm 0.811$
DINOZAURD (Ours) $\mathbf{6.96\pm 0.04}$ $\underline{2.987\pm 0.003}$ $0.272\pm 0.002$ $\underline{13.479\pm 0.044}$
DINOZAURL (Ours) $\underline{7.11\pm 0.05}$ $3.041\pm 0.008$ $\underline{0.266\pm 0.001}$ $13.749\pm 0.003$
DINOZAURB (Ours) $7.49\pm 0.07$ $\mathbf{2.767\pm 0.011}$ $\mathbf{0.168\pm 0.001}$ $\mathbf{11.667\pm 0.136}$
@@ 4.2 Results: uncertainty quantification
In this section, we assess the UQ performance of the Bayesian version of our model, DINOZAURB, introduced in Section 3.3, on the more challenging non-uniform data. As baselines, we consider MC Dropout [16] (NOD) and Laplace approximation [44] (NOL). MC Dropout is a widely used, model-agnostic UQ method, which we apply by enabling 10% dropout in the Fourier blocks’ MLPs to emulate stochasticity in the message passing. The Laplace approximation follows [44], modeling uncertainty via a fitted distribution over the parameters of the final MLP layer $\mathcal{Q}$. The discussion of hyperparameter choice, including $\mu_{\text{prior}},\sigma_{\text{prior}}$ required for setting $\ln\tau$ distributions is in the Appendix B.
UQ metrics in Table 2 are calculated by drawing 100 posterior predictive distribution samples per test data instance. The reported mean and standard deviation are obtained from three independently trained models with random weight initialization. Metrics are computed using the Uncertainty Toolbox [61]; full metric definitions are provided in the Appendix D. Aleatoric noise was added at prediction time to all model outputs, using a fixed $\sigma^{2}=e^{-4}$ for dropout and Laplace models, and a learned variance term for Bayesian DINOZAUR that converged to $\sigma^{2}\approx e^{-4.8}$. DINOZAURB matches or surpasses all baselines, achieving lower negative log-likelihood (NLL) than dropout models or Laplace approximation. It also outperforms all baselines in terms of miscalibration area (MA) and interval score (IS). An important consideration is that the Laplace approximation requires a secondary computationally expensive step to obtain the Hessian of the loss function with respect to the network parameters after training. In contrast, DINOZAURB natively supports UQ, thereby providing accurate uncertainty estimates out of the box. See results in Figure 3.
@@ 4.3 Ablation studies
To evaluate the scalability of DINOZAUR, we conduct experiments varying model width, depth, and the number of truncation modes using the ShapeNet car dataset. As shown in Figure 4(a), test error (measured by RL2 loss) decreases with increasing architectural hyperparameters, saturating near the values used in our main experiments. Further improvements may be limited by the input data resolution of $64^{3}$. Figure 4(b) confirms our scaling claims: depth increases parameter count almost linearly, width affects it quadratically, while modes do not affect the model size.
@@ 5 Discussion
In our investigations, we found that the primary limitation of the diffusion mechanism is its radial symmetry. We observed that diffusion alone struggles to train well. Our solution is to include gradient features, though at the cost of increased computational time. Investigating other strategies, like introducing anisotropy intrinsically, using the modifications proposed in [62], or mixing gradients with other features before applying inverse FFT overhead, are promising directions for future work. We found that our Bayesian formulation is sensitive to hyperparameter settings of the prior distributions. Customizing priors across network layers to capture different physical properties could improve the robustness and interpretability of the model. More broadly, exploring other parsimonious kernel parameterizations may offer both interpretability and improved generalization, especially in settings with limited data or strong physical priors. Finally, since our parameterization is independent of spatial dimensionality, it opens up opportunities for pretraining on lower-dimensional or simpler physical systems, with the potential to accelerate learning in more complex, higher-dimensional domains through transfer learning [63].
@@
Acknowledgments and Disclosure of Funding
We want to thank Greg Bellchambers, Bachir Djermani, Axen Georget, Pavel Shmakov, and Phoenix Tse for insightful discussions and collaboration on the implementation.
CRediT author statement: Albert Matveev: Conceptualization, Methodology, Software, Investigation, Writing - Original Draft, Visualization. Sanmitra Ghosh: Conceptualization, Methodology, Software, Investigation, Writing - Review and Editing, Visualization. Aamal Hussain: Software, Investigation, Writing - Original Draft. James-Michael Leahy: Conceptualization, Methodology, Software, Writing - Review and Editing. Michalis Michaelides: Conceptualization, Resources, Writing - Original Draft, Supervision, Project administration.
@@ Appendix A Benchmarks
Darcy flow 2 Regular $[241;241]$ Regular $[241;241]$ 1024, 1024
NS 2D+T 2+1 Regular $[64;64;10]$ Regular $[64;64;10]$ 1000, 200
ShapeNet car 3 Regular $[64;64;64]$ Irregular [3,586][3{,}586] 561, 100
Ahmed bodies 3 Regular $[64;64;64]$ Irregular $\sim[150{,}000]$ 500, 51
@@
Darcy flow Permeability Standard $[271;271]$ Pressure Standard
@@
ShapeNet car SDF None $[72;72;72]$ Pressure Standard
Ahmed bodies SDF, Inlet velocity None $[72;72;72]$ Pressure Standard
@@
This dataset contains steady-state 2D solutions of the Darcy flow equation on the unit box, modeling the flow of fluid through a porous medium [9]. It includes samples with spatial resolution of $241\times 241$ points. The network received a scalar permeability field as input and was tasked to predict scalar pressure values. To ensure non-periodicity, we applied padding during training, increasing the resolution to $271\times 271$. A standard scaler was applied to both input and output fields. Train and test splits had 1024 samples each.
@@
The Navier–Stokes dataset includes simulations of incompressible, viscous fluid flow with constant density and viscosity $\nu=10^{-5}$ [9]. It features a periodic 2D domain of resolution $64\times 64$, with 20 time steps of the vorticity field evolution. We trained models to predict the final 10 time steps from the initial 10, using 1000 training and 200 test samples. Inputs were composed as 3D tensors containing 10 time steps for every spatial point. Padding was introduced only to the time dimension, making the input to be of $64\times 64\times 15$. We appended spatial coordinates uniformly sampled in $0,2\pi)$ and a time coordinate with an integer index $t=[0,1,\dots 9]$, which resulted in 4 input channels that were mapped to one channel.
@@
This benchmark contains 661 car meshes paired with Reynolds-Averaged Navier–Stokes (RANS) simulations [59]. We focus on predicting the scalar pressure field at the mesh vertices, which are fixed in size at 3,586 points per mesh. The dataset is split into 561 training and 100 test samples. Target pressures were standardized during training. Following Li et al. [10], inputs are constructed by sampling a fixed-resolution $64^{3}$ grid over the global bounding box encompassing all meshes. A signed distance function (SDF) is computed per mesh. Grids are padded to $72^{3}$ before being passed to the network.
@@
The Ahmed bodies dataset [60] follows a similar setup to the ShapeNet car but introduces varying inlet velocities as an additional signal. This constant is appended to the SDF input, resulting in two input channels. Meshes vary in size, containing between 90,000 and 200,000 vertices. The grid resolution, padding, and target scaling match those used in the ShapeNet car.
The GINO-decoder architecture [10] used for both ShapeNet and Ahmed datasets requires precomputed fixed-radius neighborhoods to map grid samples to mesh vertices. We use the radii from the original paper: $0.05$ for the ShapeNet car and $0.035$ for Ahmed bodies.
For models trained on non-uniform data, we additionally apply Transformer-style positional encoding [64] to the input features, grid locations, and mesh vertices.
@@ Appendix B Configurations of trained architectures
TGINOD 64 [24, 24, 24] $10^{-4}$ False $\mathrm{D}=0.1,\,\sigma^{2}=e^{-4}$
@@
TGINOL 64 [24, 24, 24] $10^{-4}$ False $C=500,$
DINOZAURL 64 [24, 24, 24] No True $\alpha=10^{6}$
@@
All results reported in our paper are from experiments conducted specifically for this study to ensure a fair and consistent comparison across models. Architectural settings for the baselines are summarized in Table 5. All models featured 4 FNO blocks with GELU activations. The channel width $d_{c}$ and the number of modes were selected to align with the original configurations of FNO [9] and GINO [10], with the exception of Navier-Stokes data. We increased modes and channels to reflect the complexity of this benchmark. For models employing Tucker factorization, the rank $r$ was either adopted directly from the original implementation (GINO) or chosen to match the total number of parameters in DINOZAUR.
@@
where the expectation is approximated via Monte Carlo samples, and $C$ and $\alpha$ are hyperparameters selected on a validation set. In our experiments, we set $C=500,\,\alpha=10^{6}$
@@
Optimization hyperparameters are detailed in Table 6. All models were trained using either AdamW Schedule-Free [57] or AdamW combined with a One-Cycle scheduler [56], weight decay set to $10^{-4}$. Both options allow for faster convergence with fewer epochs and reduce sensitivity to hyperparameter choices. The final setup was selected by comparing both options and reporting the results for the configuration that achieved the best performance. A limited learning rate sweep was conducted in the range $10^{-4},10^{-2}]$, with early termination applied to underperforming runs. For the One-Cycle scheduler, the learning rate in the table refers to the max LR parameter.
@@ Appendix C Distributions of diffusion times
Figure 7: Lower diagonal full-rank learned covariance matrices of the posterior distributions. Each matrix represents a $d_{c}\times d_{c}$ ($64\times 64$) covariance matrix of the multi-dimensional distribution of diffusion times within each block.
@@ Appendix E Elaboration on the universal approximation
[…]
@@
making a full network $\mathcal{N}_{\theta,\infty}=\mathcal{Q}\circ\mathcal{N}_{\theta_{1},\infty}\circ\mathcal{P}$ an averaging neural operator in the sense of Equation 2.3 in [66] on the torus $\mathbb{T}^{d}$. By Corollary 3.3 [66], $\mathcal{N}_{\theta,\infty}$ is a universal approximator on the torus. Here, we note that the lifting and projection operators defined in [66] have an explicit positional dependency, and thus, one should understand them as periodized MLPs.  [^1] Note that Theorem 2.1 in [66] as it is formulated, applies to bounded Lipschitz domains. Thus, the proof of Corollary 3.3 [66] is not immediate. One must either follow the proof of Theorem 2.1 to see that it applies to the torus, or first formulate it on the domain $(0,2\pi)^{d}$, invoke the Theorem 2.1, and then argue that it holds when one restricts to periodic functions on $(0,2\pi)^{d}$.
Now consider a general domain $\Omega\subset(0,2\pi)^{d}$. By Lemma 41 by Kovachki et al. [67] and Lemma A1 in [66], there exists a continuous, linear extension operator $\mathcal{E}:C^{s}(\Omega;\mathbb{R}^{d_{a}})\rightarrow C^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{a}})$ such that $\mathcal{E}[a]|_{\Omega}=a$ and $\mathcal{E}[a]$ is periodic on $0,2\pi]^{d}$ for all $a\in C^{s}(\Omega;\mathbb{R}^{d_{a}})$. We note that the conditions of the Theorem 9 in [67] are satisfied. By following the proof of this theorem, we apply the extension operator $\mathcal{E}$ and use the universal approximation on the torus, establishing universal approximation on $\Omega$. ∎